OpenStudy (anonymous):

assume the Pi = 101 kpa an vi= 10.0l if pf=43.0 kpa what is v

jimthompson5910 (jim_thompson5910):

Pi*Vi = Pf * Vf 101*10.0 = 43.0 * x solve for x

OpenStudy (anonymous):

how do you find out what x is ? im just learning this

jimthompson5910 (jim_thompson5910):

you divide both sides by 43.0 to isolate x

OpenStudy (anonymous):

so i just do 43.0 divided by 43.0 ?

jimthompson5910 (jim_thompson5910):

yes on the right side do the same on the left side

OpenStudy (anonymous):

um well what sides? sorry... im really confuessed because its only on a piece of paper my teacher wrote it. its only the problum

jimthompson5910 (jim_thompson5910):

101*10.0 = 43.0 * x 101*10.0/43.0 = 43.0 * x/43.0 101*10.0/43.0 = 1x x = 101*10.0/43.0 x = ???

OpenStudy (anonymous):

23.488372093?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

do u know how to do (9x10 power of -1) x 3 x 10 power of 10 )?

jimthompson5910 (jim_thompson5910):

multiply the bases (9 and 3) add the exponents (-1 and 10) then convert to scientific notation (if necessary) or standard form

OpenStudy (anonymous):

its standard form how do u write that 9x3=27 an -1 +10 =9

jimthompson5910 (jim_thompson5910):

so it would be 27 x 10^9 you can convert to scientific notation or standard form if you want

jimthompson5910 (jim_thompson5910):

the answer in scientific notation would be 2.7 x 10^10 in standard notation, it would be 27,000,000,000 (which is the number 27 billion)

OpenStudy (anonymous):

thank you i have a few more questions if u dont mind

OpenStudy (anonymous):

if electirc forces are so strong why dont i notice them acting on everything around me ? a friend asks how do i respond ? thats the question

jimthompson5910 (jim_thompson5910):

well there's a balance in forces that keeps everything from either flying apart or pulling together and collapsing in on itself

jimthompson5910 (jim_thompson5910):

you only see an imbalance in force when you see things like electricity or magnetism

OpenStudy (anonymous):

okay thank you show how multiply both sides of the acceleration equation a= fnet (over) m by m results in the net force equation fnet+ma

jimthompson5910 (jim_thompson5910):

I'm using f instead of fnet a = f/m a*m = f/m * m ... Multiply both sides by m. a*m = fm/m a*m = f*1 a*m = f f = a*m f = ma

OpenStudy (anonymous):

i dont get it ......

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

all of it i dont know what m is ? or like any of it?

jimthompson5910 (jim_thompson5910):

there's no need to know m

jimthompson5910 (jim_thompson5910):

you just need to rewrite a = f/m into f = ma

OpenStudy (anonymous):

ohh okay i get it . hiw do u do calculate the acceleration of a 25 kg object that is moved with a force of 300 n sorry i have to have this an an whole other book done in an hour

jimthompson5910 (jim_thompson5910):

f = ma 300 = 25*a solve for 'a'

OpenStudy (anonymous):

12?

jimthompson5910 (jim_thompson5910):

good, the acceleration is 12 m/s^2

OpenStudy (anonymous):

okay thanks the mass of the following box is 120.00g calculate the destiny the box has a 5cm 6 cm an a 4 cm by it ?

jimthompson5910 (jim_thompson5910):

first find the volume of the box

OpenStudy (anonymous):

120 ?

jimthompson5910 (jim_thompson5910):

now divide the mass by the volume

jimthompson5910 (jim_thompson5910):

D = M/V D = 120/120 D = 1

jimthompson5910 (jim_thompson5910):

so the density is 1 gram per cubic cm

OpenStudy (anonymous):

so is that the answer ?

OpenStudy (anonymous):

theres another question the normal states of matter are solid liquid gas a friend says then there is a really uncommon state called plasma how would you respond to this claim

jimthompson5910 (jim_thompson5910):

I would respond by showing him/her this link http://en.wikipedia.org/wiki/Plasma_%28physics%29

jimthompson5910 (jim_thompson5910):

this link shows there are tons of places where you would find plasma so it's not so rare

OpenStudy (anonymous):

okay hah thanks evaluate this statment the object is in motion in a stright line and it is moving with a constant speed that must mean that no force is acting on it

jimthompson5910 (jim_thompson5910):

ok just say whether it's true or false?

OpenStudy (anonymous):

is it false ?

jimthompson5910 (jim_thompson5910):

why is it false?

OpenStudy (anonymous):

because wouldnt sommething have to be moving it ?

jimthompson5910 (jim_thompson5910):

think of it out in outerspace away from any planet

jimthompson5910 (jim_thompson5910):

it's far from any forces and it moves in a straight line at the same speed

OpenStudy (anonymous):

well then its true

jimthompson5910 (jim_thompson5910):

nothing is speeding it up or slowing it down nothing is changing its direction

jimthompson5910 (jim_thompson5910):

so there are no forces acting on it

jimthompson5910 (jim_thompson5910):

yes it's true

OpenStudy (anonymous):

okay well how do u do this it says place a nymber of the law of motion w=in the blank next to rhe examples of the law giving the first on is fnet =ma

OpenStudy (anonymous):

2 seven ball is struck by the cue ball 3 an air filled balloon is released 4 you push on a box of books

jimthompson5910 (jim_thompson5910):

not sure what this is saying

jimthompson5910 (jim_thompson5910):

something seems missing

OpenStudy (anonymous):

ill type it again hold on

OpenStudy (anonymous):

place the number of the law of motion in the blank next to the examples of the law giving the examples are

OpenStudy (anonymous):

*fnet =ma * seven ball is struck by the cue ball * an air filled balloon is released *you push on a box of books

jimthompson5910 (jim_thompson5910):

oh i gotcha, ok one sec

jimthompson5910 (jim_thompson5910):

fnet =ma is definitely the second law of motion

jimthompson5910 (jim_thompson5910):

even ball is struck by the cue ball sounds like the third law of motion

jimthompson5910 (jim_thompson5910):

not sure about the third one but you push on a box of books also sounds like the third law of motion

OpenStudy (anonymous):

thank you! how many cubic centemeters dose 1.304L have

OpenStudy (anonymous):

the question is the distance from the launch site to the observation is 120 meters calculate the adverage height a if the obtained altimeter reading from 4 observations sites were the following

jimthompson5910 (jim_thompson5910):

something seems missing here

jimthompson5910 (jim_thompson5910):

oh and the pic is mirrored lol can you fix that please?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u see the shpae ill typer the answers

jimthompson5910 (jim_thompson5910):

yeah i'll just try to reverse it in my head

OpenStudy (anonymous):

* 36.5 *35.0 *40.0 *38.0

OpenStudy (anonymous):

question is

OpenStudy (anonymous):

the distance from the launch site to the observer is 120 meters calculate the average height a , if the obtained altimeter readings from 4 observation sites were the following

OpenStudy (anonymous):

do u got it?

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

does it give you answer choices?

OpenStudy (anonymous):

yes its 36.5 , 35.0 40.0 or 38.0

jimthompson5910 (jim_thompson5910):

those are the heights though right?

OpenStudy (anonymous):

well it says derees after the numbes

jimthompson5910 (jim_thompson5910):

oh nvm, ok one sec

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

(120*tan(36.5) + 120*tan(35) + 120*tan(40) + 120*tan(38) )/4 = 91.8166161276659 which is roughly 91.82 meters

OpenStudy (anonymous):

so which one is it ?

jimthompson5910 (jim_thompson5910):

91.82 if you're rounding to two decimal places

OpenStudy (anonymous):

ohhh gotcha

OpenStudy (anonymous):

what bout this

OpenStudy (anonymous):

if a current flowing through a light buld is 0.75 a an the voltage diffrence is across the light bulb is 120v how much resistance does the light bulb have

jimthompson5910 (jim_thompson5910):

V = I*R 120 = 0.75*x solve for x

OpenStudy (anonymous):

61 ?

OpenStudy (anonymous):

make a general statement about the facts that increases resistance

jimthompson5910 (jim_thompson5910):

120 = 0.75*x x = 120/0.75 x = ???

OpenStudy (anonymous):

isnt it 61 ?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

then idk

jimthompson5910 (jim_thompson5910):

120/0.75 = 160